# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v1

@@ 12 Related work
Previous research has looked at the security of implanted medical devices in general [29, 16, 54], in addition to looking at insulin pumps in particular [39, 50], with more recent work looking at providing improved security [44, 11]. Also, recent work has looked at applying formal methods to insulin pumps for high assurance [49]. These works focus on the device and their communication channel. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
Our work on coping with malicious pump drivers builds on previous work on malicious device drivers in operating systems [62], [15], [27], [28], [43], [31]. Our work applies these principles to a new domain: automated insulin delivery systems.
@@ 13 Conclusion
Our audacious long-term goal is to turn a T1D diagnosis from a death sentence into an indicator of longevity, where people living with T1D will be expected to live longer than their healthy peers. This longevity will come by virtue of the tight control that they maintain over their metabolism through advanced computer systems. Our first step towards this goal is to ensure that people can use trustworthy computer systems to manage their glucose.
Appendix A Protocol for real world use
One individual, who we will call Bob, used GlucOS for 2.5 months. We put in place a protocol where we would stop our study and ask Bob to cease using our system and revert back to his previous insulin regime. In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS. We would have stopped our study if any of the following conditions occurred:
  • Excessive hypoglycemia. We define excessive hypoglycemia as having spent 5% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl. These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for Bob.
  • Insufficient CGM data. If Bob would have had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for Bob.
@@ Appendix A Protocol for real world use
  • Serious complications due to diabetes. Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where Bob was unable to recover from hypoglycemia himself and had to get help from someone else, diabetic ketoacidosis, or seizure.
  • New medical diagnosis requiring significant attention. If Bob had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for Bob.
  • New mental health diagnosis. If Bob were struggling with mental health during the study, the study would have the potential to be a distraction.
@@ Appendix B List of pump setup and UX communication messages
Appendix C CGM state machine
Inability to retrieve readings from a CGM also count as runtime failures in addition to those discussed in Section 7. However, these failures are easier to handle. In this section, we describe the CGM state machine that handles such failures. We note that transitions from the state machine described in Section 7 supersede state transitions from the CGM state machine.
@@ Appendix C CGM state machine
Appendix D Formal verification: invariants and bugs

### v2

@@ 10 Related work
In addition to the work already discussed in this paper, there are several other related works.
Previous research has looked at the security of implanted medical devices in general [32, 16, 60], in addition to looking at insulin pumps in particular [39, 56], with more recent work looking at providing improved security [45, 9]. Also, recent work has looked at applying formal methods to insulin pumps for high assurance [55]. These works focus on devices and their communication channels. In contrast, with GlucOS we assume that these devices are correct and secure and focus our efforts on the software we use to run the automated insulin delivery system.
Researchers have started to look at combining untrustworthy ML with more traditional forms of automated decision making, like Markov Decision Processes, from a theoretical perspective [41, 50, 57, 20, 40]. These theoretical works focus on providing algorithms to prove desirable properties about a combined ML / Markov system, like robustness and stability. In contrast, we take advantage of domain knowledge and formulate a simple and practical system for one specific problem: automated insulin delivery.
@@ 11 Conclusion
The design of GlucOS, while grounded in established OS and security principles, incorporates several non-obvious elements that emerged from our deep engagement with the problem domain and real-world testing.
First, our approach to algorithmic security used classic security techniques, like reference monitors and the Simplex architecture. However, it was our novel insight around all correct algorithms dosing the same amount of insulin that led us to these classic techniques. Plus, from our exploration, the most interesting part was not the ML, it was around the principles of reactive safe models and in particular our principle that these safety-critical algorithms should be reactive.
Second, our pump driver security went beyond traditional specification and checking. We needed to actively program the pump to handle dropped cancel commands, which was an unexpected but crucial insight.
Third, the biological invariant emerged as a key check in our system, addressing a blind spot in typical simulations that ignore changing human dynamics. This forced us to develop novel formal methods to handle scenarios where our glucose metabolism models could be drastically wrong.
Fourth, our approach to human interaction in the system was nuanced. We found that simply alerting humans or suspending insulin delivery were not always appropriate. For security alerts, personalization and customization were important. Stopping insulin delivery could be dangerous if the user has high glucose and insufficient insulin on board, which required more context-aware responses.
These non-obvious design decisions underscore the complexity of creating a truly trustworthy automated insulin delivery system. They reflect the value of combining theoretical security principles with practical, real-world testing and domain-specific knowledge. This interdisciplinary research drew on deep expertise from computer security, medicine, and ethics to create a practical system.
Our work improves the security of automated insulin delivery systems and also provides a framework for thinking about security in other critical healthcare technologies. Our exploration is an extreme example of personalized, distributed, and automated healthcare using trustworthy computer systems.
Appendix A Ethical and safety considerations
One individual, who we will call Bob, used GlucOS for 9 months. Six university students also used the predictive alerts from GlucOS for 2 months. We put in place a protocol where we would stop our study and ask people to stop using our system and revert back to their previous management regime. In our protocol, we have formal in-person meetings with Bob every week to check his data and make sure that he is not incurring risk as a result of using GlucOS, and weekly surveys for the six students, plus monthly meetings. We would have stopped our study if any of the following conditions occurred:
  • Excessive hypoglycemia. We define excessive hypoglycemia as having spent 8% or more of time during a week with CGM readings below 70 mg/dl or more than 1% below 54 mg/dl. These amounts of time spent in hypoglycemia would have represented a significant increase in time spent in hypoglycemia for participants.
  • Insufficient CGM data. If they had less than 70% of the time during a week with CGM readings, then the lack of CGM data would represent a decrease for the participants.
@@ Appendix A Ethical and safety considerations
  • Serious complications due to diabetes. Serious complications include any hospitalization for diabetes related issues, severe hypoglycemia where a participant was unable to recover from hypoglycemia themselves and had to get help from someone else, diabetic ketoacidosis, or seizure.
  • New medical diagnosis requiring significant attention. If anyone had received a new medical diagnosis during the study, the study would have had the potential to be a distraction for participants.
  • New mental health diagnosis. If anyone were struggling with mental health during the study, the study would have the potential to be a distraction.
In addition to our safety protocol, we followed best practices for ethical research to ensure that we conducted our study ethically.
  • IRB. Our study went through our university’s IRB process, where the study including six university students who used GlucOS’s alerting software was approved, and our case study on Bob was determined to be “self evaluation” and exempt from IRB.
  • Informed consent. All participants went through a written informed consent process.
  • Data management. GlucOS does not store any personally identifiable information and uses access control on our Google Cloud Platform infrastructure to ensure that only the authors of the paper can access this anonymized data.
  • Risk/benefit analysis. All participants are already living with T1D and use CGMs and inject insulin, so they already take on the risks inherent in managing T1D. The benefits include improved control over their T1D. There are risks in using experimental software, but all participants continued to rely on their CGM alerts as a last line of defense.
  • Participant selection. We selected participants who were university students living with T1D.
  • Adverse event reporting. Had any participants experienced adverse events during the study, we would have reported it to our IRB for guidance and addressed it according to our safety protocol.
  • Communication. All surveys were done through our university’s medical school infrastructure for private communication with individuals.
Since Bob used the full GlucOS software, including for insulin delivery, we included several other safety mechanisms:
  • Bob increased the frequency of his visits to his Endocrinologist while using the GlucOS software. He went from yearly appointments to quarterly appointments.
  • During the study, twice Bob got lab work based on blood tests to provide an independent measure of his overall metabolic health. This lab work helped to ensure that he was in good health and that there were no hidden negative impacts from his use of GlucOS.
Overall, GlucOS had a large positive impact on the individuals who used it. Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%. That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower. All of the students decreased the amount of time spent in hypoglycemia. These results are profound for the seven people using GlucOS, and it had a positive impact on their health and quality of life.
Finally, Bob was a willing and enthusiastic user of the GlucOS system. To be clear, Bob is not a graduate student who was forced to use GlucOS so that his advisor could publish a paper. Rather, Bob struggled with the cognitive load of T1D management and wanted an automated insulin delivery system to offload part of the burden. However, he was unwilling to use any of the other automated insulin delivery systems out there due to concerns over their security. Bob found that GlucOS, with its focus on security, correctness, and simplicity, met his needs.
@@ Appendix B List of pump setup and UX communication messages
Appendix C Runtime failure management
GlucOS’s state machine consists of three operating states: (1) ML with safety operating in closed loop, (2) reactive safe model operating in closed loop and (3) manual dosing in closed loop. ML provides the tightest level of control in managing glucose but poses the largest immediate threat from excess insulin since it preemptively injects insulin by anticipating rise in glucose. The reactive safe model provides slightly lesser control, but poses milder immediate threat since it only injects insulin to account for measured excess glucose. Manual dosing offers the least amount of control but poses no threat since it only injects the minimum insulin needed for sustenance. GlucOS predominantly operates in the state running ML (with safety) in closed loop.
Inability to retrieve readings from a CGM also count as runtime failures. However, these failures are easier to handle.
@@ Appendix C Runtime failure management
Appendix D Did formal verification help with correctness?
Appendix E Formal verification: invariants and bugs